The sum of three consevutive odd intergers is 195? What are the intergers?
1st?
2nd?
3rd?

Found 2 solutions by mananth, amoresroy:
Answer by mananth(16946)

let the first integer be x
x+x+2+x+4=195
3x+6= 195
3x=189
x= 63
63, 65,67

Answer by amoresroy(361)

The sum of three consevutive odd intergers is 195? What are the intergers?
Let x = the smallest integer
x + (x+2) + (x+4) = 195
3x = 195-6
3x = 189
x = 63
The integers are 63, 65, and 67.